OpenStudy (anonymous):

A person of mass 60 kg is inside a lift of mass 940 kg and presses the button on control panel. The lift starts moving upwards with an acceleration 1.0 m/s2. If g510 ms22, calculate the tension in the supporting cable.

OpenStudy (ujjwal):

Are you sure you typed it correctly ... " If g510 ms22, " ? what does it mean?

OpenStudy (anonymous):

its T - 1000(g) = 1000(1) SO T=1800N

OpenStudy (anonymous):

You can consider an object with mass of 1000 kg instead of those two ones.  As shown above the \[F _{2}\] force is reaction force equal to that of gravity by magnitude and opposite in direction, and it has positive sign due to we have chosen positive direction looking upwards, the same positive sign has the force \[F _{1}\] with which some creature is pulling our lift. Let's now consider some point A on the cable and see what forces act on it:  We will see that the same two forces acts on this point of cable shown as sum of forces \[F _{total}\] and by the third law of Newton there must be another force compensating this action by its reaction, and that will be a tension force T which has negative sign in this case, but we are not interested in its direction( because it acts in either of the two directions, convince yourself by choosing a higher point on the cable) but only in magnitude, so \[T = m(a+g)=1000 kg \times (1 m/s ^{2}+10m/s ^{2})=1.10 \times 10^{4} N\]